Qd=100-5p and Qs= -20+3p if the price is R10


1
Expert's answer
2022-04-05T19:19:36-0400

"Q_d=100-5p,"

"Q_s=-20+3p,"

"Q_d=Q_s,"

"100-5p=-20+3p,"

"120=8p,"

"p=15,"

"Q_d=Q_s=25."